OpenStudy (anonymous):

We often say that log (1+x) = x - x^2/2 + x^3/3 etc. Is the statement true for x = 500 (or may be larger values?) If not, why? Can anyone give a definite answer to this?

ganeshie8 (ganeshie8):

@eliassaab

OpenStudy (anonymous):

The above series converges for |x|<1 So the answer is no for x=500

OpenStudy (anonymous):

\[ \ln (1+500.)\approx6.21661\\ \frac{x^3}{3}-\frac{x^2}{2}+x=4.15422\times 10^7 \text{ fpr } x=500 \]

OpenStudy (anonymous):

So,due to convergence it is not possible. But, can u say for what values the series is possible.

OpenStudy (tkhunny):

Series Convergence is NOT a finite process. Any time you cut it off at some finite point, there is no equality. 0) Equality - Not Finite However, with appropriate error estimates, you may be able to (in decreasing order of precision): 1) Get close enough so that the error cannot be measured by any existing device. 2) Get close enough to build a useful device. 3) Get close enough to verify homework results.